4.00×10−3 molmol of HBrHBr are dissolved in water to make 11.0 LL of solution. What is the concentration of hydroxide ions, [OH−][OH−], in this solution?


1
Expert's answer
2021-11-15T10:58:26-0500

4.004×10-3/11= 3.64×10-4M

OH- = 10-14

10-14/3.64×10-4M= 2.75×10-11